A man puts $10 in the bank for his son on each of his birthdays from the first to the twentieth inclusive. If the money accumulates at 3% compound interest, what is the total value on his son's twenty first birthday?
1
Expert's answer
2014-04-23T12:04:09-0400
Future value is calculated by the following formula:

FV=PV(1+i)^n=10(1+0.03)^20=10*1.03^20≈18.06,

where PV=$10 is the present value, i=0.03 is the effective interest rate per year, n=20 is the number of periods.

Answer: $18.06